# Satori: Towards Proactive AR Assistant with Belief-Desire-Intention User Modeling

### v2

@@ Abstract.
Augmented Reality assistance are increasingly popular for supporting users with tasks like assembly and cooking. However, current practice typically provide reactive responses initialized from user requests, lacking consideration of rich contextual and user-specific information. To address this limitation, we propose a novel AR assistance system, Satori, that models both user states and environmental contexts to deliver proactive guidance. Our system combines the Belief-Desire-Intention (BDI) model with a state-of-the-art multi-modal large language model (LLM) to infer contextually appropriate guidance. The design is informed by two formative studies involving twelve experts. A sixteen within-subject study find that Satori achieves performance comparable to an designer-created Wizard-of-Oz (WoZ) system without relying on manual configurations or heuristics, thereby enhancing generalizability, reusability and opening up new possibilities for AR assistance.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXsubmissionid: 4490ccs: Human-centered computing Mixed / augmented reality
Figure 1. Satori is a mind-reading monkey-shaped monster in Japanese folklore. We named our system Satori to highlight the importance of understanding the user state (human mind) while building proactive AR assistants. The Satori system combines the users’ self-knowledge and task goal of the immediate user action with a large-language model (LLM) model to provide visual assistance relevant to the user’s immediate needs. This kind of proactive AR assistance is achieved by implementing the Belief-Desire-and-Intention model based on two formative studies with 12 experts. Here the belief component reflects whether the users know where the task object is, and how to perform certain tasks (e.g., knowledge level); the desire component is the actionable goal; and the intent component is the immediate next step needed to complete the actionable goal. Our code will be open-sourced upon acceptance.
@@ 1. Introduction
Building on the findings from the first study, the second formative study engaged six experts—three human-computer interaction (HCI) researchers and three psychology researchers—in dyadic interviews to explore design strategies for a more proactive AR assistance. The design sessions found four key design considerations: 1) understanding human actions; 2) recognizing surrounding objects and tools; 3) assessing the current task; and 4) anticipating immediate next steps. Following the suggestions from the experts, these findings were later integrated with the well-established belief-desire-intention (BDI) model (Georgeff et al., 1999; Leslie et al., 2005; Bratman, 1987; Cohen and Levesque, 1990), resulting in an AR-specific adaptation that guided the development of our system, Satori.
@@
    A 16-user empirical study showing that Satori delivers performance comparable to designer-created AR assistance in terms of timing, comprehensibility, usefulness, and efficacy.
@@ 2. Related Work
Pre-designed assistance in AR and VR applications typically involves preset rules for actions or reminders triggered by specific user inputs or situations. This kind of rule-based assistance is simple and intuitive, providing users with readily available support that can be accessed on demand or in time sequences (Lapointe et al., 2022). While implementing such assistant systems is straightforward (Sara et al., 2022), they require extensive manual user interaction to describe and confirm the user’s needs. For example, Sara et al. demonstrated an AR maintenance assistant wherein the technician needed to manually confirm the completion of each step and proceed to the next step using touchpad controls or voice commands (Sara et al., 2022).
@@ 3.2. Tasks
In the first session, we collected participants’ prior working experience using AR assistants, the challenges they faced in creating them, and their assessment of the assistants’ potential benefits and applications. Additionally, we discussed the concept of proactive AR assistance with participants and collected their insights on potential benefits and use scenarios. In the second session, participants were asked to design AR assistants for two everyday scenarios out of the four. These two scenarios were assigned in a pre-determined order to balance the total number of designs. We use WikiHow  [^1] https://www.wikihow.com/ to obtain detailed, step-by-step instructions as the task background information for participants. These instructions ($averagesteps:\bar{x}=7$) provide the framework to make guidance, and participants can elaborate (e.g., adding additional steps) at their will. Aside from the text instructions, we recorded videos in first-person view using the original instructions to provide visual reference and interaction context for participants. Given instructions, images, and videos depicting the scenarios, participants were asked to design: 1) if a piece of guidance is needed for a particular step; 2) when the guidance should appear and for how long; 3) the modality of the guidance; 4) the content of guidance. The above questions focus on the questions of “if”, “when”, “how”, and “what” in AR assistance, which is a common architecture for guiding users in the literature and current practice (Lindlbauer et al., 2019).
@@ 3.4.4. Challenges of Proactive AR Assistance
Four participants (P2, P4, P5, and P6) emphasized the difficulty of accurately detecting users’ intentions in AR. P5 brought up the limited field of view (FoV) in AR headsets and the low accuracy of detection algorithms as two main issues, although the former (limited FoV) might be amongst the causes of the latter (low accuracy). P5 commented that “…sometimes, the system might trigger guidance when the user doesn’t need it, which could lead to confusion…” Similarly, P4 discussed how AR software in the industry has struggled to fully apprehend complex user environments and actions, causing confusion. This view is also shared by P2, who mentioned that proactive assistance might confuse users if it lacks self-explanatory features. P2 stated that “if (the system is) fully automatic, you need the system to have some type of feedback. Automation without feedback may confuse the user.
@@ 3.4.5. Design results for four common scenarios
Table 1. Types of assistance provided across different modalities suggested by AR design experts. The overlays are used to indicate locations or to indicate how to interact with apparatus in the scene; a progress indicator reflects how far the user is into the task. The image and animation are designed to provide illustration of actions, positions, and to show “how” to complete the current step; The checkpoint cue, according to participants, are used to indicate step completion. The timer is designed to count time for time-sensitive steps such as making pour-over coffee.
@@ 4.2. Known Challenges
Table 2. Expert background in the co-design. We paired one computer science expert with one psychology expert per each group. In total three groups participated the co-design.
@@
(a) Initially presented diagram and the available modules. (b) Sample result from Group B.
Figure 2. Participatory design in the first session. The experts need to collaborate on creating a desired assistant framework based on the provided baseline diagram. At the bottom of Figure (a), the experts can find the system components for perception.
@@ 4.6. Results
When describing ideas to implement the BDI model within the AR context, the expert groups outlined how belief supports the filtering of duplicated or unnecessary assistance and acts as a screening step to narrow the assistance’s scope. They further outlined that desire refers to the goals of a given task. In AR, this means the system should model the user’s actions and goals (Group B and C). Finally, the expert groups indicated that intention comprises a small step toward the goal and affects the timing and content of the AR assistance (Group A and C). Together these adaptations of the BDI model help to construct a novel pipeline toward proactive AR assistance.
@@
Table 3. The table summarizes main results from two formative studies.
@@ 5. Design Requirements
    Assistance should try to be transparent to gain users’ trust, feeding back the system’s reasoning and detection, and provide easily accessible information about current and overall goals in the AR environment.
@@ 6.1. Implementing the BDI Model for AR Assistance
Table 4. This table illustrates how three components— Belief, Desire, and Intention– in the BDI model are adapted for AR task guidance. BDI Comp. refers to the BDI Component; and AR Guid. Comp. refers to AR Task Guidance Component. Belief is represented through scene understanding, task-relevant object detection, and user action history to minimize distractions, locate objects, and avoid repeated instructions. Desire captures the user’s high-level task goals, inferred through LLM-based scene analysis and confirmed by the user to ensure accuracy. Intention includes predicting the next intended action using GPT-based inference with chain-of-thought reasoning and determining the timing of next actions with checkpoint-based early forecasting, reducing latency in guidance delivery.
@@
Figure 3. Overview of BDI user model. The system processes inputs from the camera’s view, dialogue (voice communication between the user and the GPT model), and the historical logger (records of prior assistance). These inputs are analyzed to infer the user’s Belief, Desire, and Intention (BDI). The inferred BDI states are utilized by the LLM to generate proactive guidance and determine the appropriate modality for delivery. The system incorporates a scene detection model to identify distractions, filtering out distracted frames to maintain stability. A task planner LLM generates a step-by-step task plan based on the inferred desire, with multiple checkpoints assigned to each step. These checkpoints are monitored by the action finish detection module, which determines task completion by verifying checkpoint progress. To minimize latency, the system employs an early forecasting mechanism that continuously evaluates checkpoint completion. Once an action is confirmed as finished, the system schedules guidance timing for the next step.
@@ 6.5. Task Planner for Checkpoints
Figure 4. Comparison of naively generated image from GPT model (i.e., Raw) with our proposed prompt (i.e., Ours). (a) “One hand presses a white button on a white espresso machine. A large red arrow points to the button. No background, in the style of flat, instructional illustrations. Accurate, concise, comfortable color style.” (b) “One hand presses a white button on a white espresso machine.” (c) “Cut stem of a red flower up from bottom, with white scissors at 45 degrees. One big red arrow pointing to bottom of the flower stem. In the style of flat, instructional illustrations. No background. Accurate, concise, comfortable color style.” (d) “Cut stem of a red flower up from bottom with white scissors at 45 degrees.
@@ 7.1. Conditions
Figure 6. Evaluation tasks using either Satori or a Wizard-of-Oz baseline. (a) The participant is assembling a mop during the task room cleaning; and (b) The participant is connecting a HDMI cable to a Nintendo Switch dock during the task connecting Nintendo Switch; and (c) The participant is preparing a filter during the task of making coffee; and (d) The participant is trimming flower stems during the task of arranging flowers.
@@ 8.1. System Evaluation
Task Action Recognition Desired Task Recognition Action Recognition Desired Task Recognition
Li et al. (Li et al., 2015) 62.10 N/A N/A N/A
Satori 79.50 100.00 85.00 100.00
Table 5. Table showing the results for the desired task and action recognition. We evaluated Satori on the GTEA dataset (Li et al., 2015) and our dataset over the top model on GTEA’s leaderboard.
@@
Task QA Highlight Object Understanding
HCRN (Le et al., 2021) 69.50 N/A
@@
Table 6. Belief inference includes the scene’s objects, location, and its interaction states with users. We evaluated this module using the EgoTaskQA dataset and our dataset to compare with the HCRN model. The EgoTaskQA dataset consists of questions about humans’ beliefs in the world and the model’s understanding of humans’ beliefs. For our dataset, three experimenters annotated the highlighted object labels, locations, and interaction states separately. Satori reached a similar accuracy ($66.50\%$) to that of the HCRN model ($69.50\%$).
Look-Ahead Time GTEA Our Dataset
Recall Precision F1 Recall Precision F1
@@
Table 7. Module-level evaluation on intention (action) forecast. We evaluate our methods on the GTEA dataset and our dataset. Three experimenters annotated user action labels in our dataset. Aside from the settings Satori use (Look-Ahead Time = 0s), we also present results for two other hypothetical conditions if we predict the action 1s or 3s earlier. For our settings, our methods reached $78.38\%$ on GTEA dataset and $62.52\%$ on our dataset.
@@ 8.2. Usability Scale
Table 9. Results for the Usability Scale and Its Non-Inferiority Tests. The table summarizes the mean scores and 95% confidence intervals (CI) for each system (our Satori system and WoZ designed by AR designer) across various user experience questions related to system usability and cognitive load. The “Vanilla” columns provide the Wilcoxon signed-rank test results (W statistic and p-values) for significant differences between systems. The “Non-Inferiority” columns show W statistics and p-values testing if Satori’s performance is non-inferior to WoZ within a set margin. Green-highlighted cells indicate established non-inferiority, suggesting that Satori performs as well as or better than WoZ in user guidance and cognitive demand.
Figure 7. User ratings on the system usability using a seven-point Likert scale. There are twelve participants in the study and the ratings are encoded by the color from red to blue. The figure compares the responses for Satori and WoZ  systems across four tasks: Arranging Flowers, Making Coffee, Cleaning the Room, and Connecting a Console. Each bar represents the distribution of responses for a specific usability question, highlighting differences in user satisfaction, comprehensibility, and task support provided by both systems.
@@
Table 10. Results for NASA TLX Questions and Non-Inferiority Tests. This table shows the mean scores and 95% confidence intervals (CI) for Satori and WoZ systems across six dimensions: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. The Vanilla Wilcoxon signed-rank test results and non-inferiority tests (highlighted in green) indicate whether the Satori system performs comparably or better than the WoZ system in terms of cognitive load.
Figure 8. NASA TLX box plot for WoZ  and Satori. The plot illustrates the distribution of cognitive load ratings across six dimensions: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Each box represents the interquartile range (IQR) with the median marked by a horizontal line, showing the variability and central tendency of participants’ workload ratings for both systems. The comparison highlights differences in perceived workload between the WoZ   and Satori conditions, providing insights into the effectiveness and usability of each approach.

### v3

@@ Abstract.
Augmented Reality (AR) assistance is increasingly used for supporting users with physical tasks like assembly and cooking. However, most systems rely on reactive responses triggered by user input, overlooking rich contextual and user-specific information. To address this, we present Satori, a novel AR system that proactively guides users by modeling both – their mental states and environmental contexts. Satori integrates the Belief-Desire-Intention (BDI) framework with the state-of-the-art multi-modal large language model (LLM) to deliver contextually appropriate guidance. Our system is designed based on two formative studies involving twelve experts. We evaluated the system with a sixteen within-subject study and found that Satori matches the performance of designer-created Wizard-of-Oz (WoZ) systems, without manual configurations or heuristics, thereby improving generalizability, reusability, and expanding the potential of AR assistance. Code is available at https://github.com/VIDA-NYU/satori-assistance.
@@
journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714188isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing User models
Figure 1. Satoriis a mind-reading monkey-shaped creature in Japanese folklore. Our system extends this concept to highlight the importance of incorporating the user’s state (i.e., knowledge and intentions) while building proactive AR assistants. The Satori system combines the tracked objects, the surrounding environment, task goals, and user actions with a large-language model (LLM) model to provide AR assistance to the user’s immediate needs. This kind of proactive AR assistance is achieved by implementing the Belief-Desire-and-Intention (BDI) psychological model with advice from two formative studies with a total of 12 experts. The belief component reflects whether the users know where the task object is, and how to perform certain tasks (e.g., task goals, high-level knowledge); the desire component is the actionable goal; and the intention component is the immediate next step needed to complete the actionable goal. Our code is available at https://github.com/VIDA-NYU/satori-assistance.
@@ 1. Introduction
Building on the findings from the first study, the second formative study engaged six experts—three human-computer interaction (HCI) researchers and three psychology researchers—in dyadic interviews to explore design strategies for a more proactive AR assistance. The design sessions found four key design considerations: 1) understanding human actions; 2) recognizing surrounding objects and tools; 3) assessing the current task; and 4) anticipating immediate next steps. Following experts’ suggestions, these findings were later integrated with the well-established belief-desire-intention (BDI) model (Georgeff et al., 1999; Leslie et al., 2005; Bratman, 1987; Cohen and Levesque, 1990), resulting in an AR-specific adaptation that guided the development of our system, Satori.
@@
    A 16-user empirical study shows that Satori delivers performance comparable to designer-created AR assistance in terms of timing, comprehensibility, usefulness, and efficacy.
@@ 2. Related Work
Predesigned assistance in AR and VR applications typically involves preset rules for actions or reminders triggered by specific user inputs or situations. This kind of rule-based assistance is simple and intuitive, providing users with readily available support that can be accessed on demand or in time sequences (Lapointe et al., 2022). While implementing such assistant systems is straightforward (Sara et al., 2022), they require extensive manual user interaction to describe and confirm the user’s needs. For example, Sara et al. demonstrated an AR maintenance assistant wherein the technician needed to manually confirm the completion of each step and proceed to the next step using touchpad controls or voice commands (Sara et al., 2022).
@@ 3.2. Tasks
In the first session, we collected participants’ prior working experience using AR assistants, the challenges they faced in creating them, and their assessment of the assistants’ potential benefits and applications. Additionally, we discussed the concept of proactive AR assistance with participants and collected their insights on potential benefits and use scenarios. In the second session, participants were asked to design AR assistants for two everyday scenarios out of the four. These two scenarios were assigned in a pre-determined order to balance the total number of designs. We use WikiHow  [^1] https://www.wikihow.com/ to obtain detailed, step-by-step instructions as the task background information for participants. These instructions ($averagesteps:\bar{x}=7$) provide the framework to make guidance, and participants can elaborate (e.g., adding additional steps) at their will. Aside from the text instructions, we recorded videos in first-person view using the original instructions to provide visual reference and interaction context for participants. Given instructions, images, and videos depicting the scenarios, participants were asked to design: 1) if a piece of guidance is needed for a particular step; 2) when the guidance should appear and for how long; 3) the modality of the guidance; 4) the content of the guidance. The above questions focus on the questions of “if”, “when”, “how”, and “what” in AR assistance, which is a common architecture for guiding users in the literature and current practice (Lindlbauer et al., 2019).
@@ 3.4.4. Challenges of Proactive AR Assistance
Four participants (P2, P4, P5, and P6) emphasized the difficulty of accurately detecting users’ intentions in AR. P5 brought up the limited field of view (FoV) in AR headsets and the low accuracy of detection algorithms as two main issues, although the former (limited FoV) might be among the causes of the latter (low accuracy). P5 commented that “…sometimes, the system might trigger guidance when the user doesn’t need it, which could lead to confusion…” Similarly, P4 discussed how AR software in the industry has struggled to fully apprehend complex user environments and actions, causing confusion. This view is also shared by P2, who mentioned that proactive assistance might confuse users if it lacks self-explanatory features. P2 stated that “if (the system is) fully automatic, you need the system to have some type of feedback. Automation without feedback may confuse the user.
@@ 3.4.5. Design results for four common scenarios
Table 1. Types of assistance provided across different modalities suggested by expert AR designers. The overlays are used to indicate locations or to indicate how to interact with apparatus in the scene; a progress indicator reflects how far the user is into the task. The image and animation are designed to illustrate actions and positions and show “how” to complete the current step. The checkpoint cues, according to participants, are used to indicate step completion. The timer counts time for time-sensitive steps, such as making pour-over coffee.
@@ 4.2. Known Challenges
Table 2. The table shows the experts’ backgrounds in the co-design. We paired one computer science expert with one psychology expert per group. In total, three groups participated in the co-design.
@@
(a) Initially presented diagram and available modules. (b) Sample results from Group B.
Figure 2. During the first session (participatory design), experts need to collaborate on creating an ideal assistant framework based on the presented diagram and modules. At the bottom of Figure (a), the experts can find the system components for perception. Figure (b) is a result of the original diagram illustrated by one expert group.
@@ 4.6. Results
When describing ideas to implement the BDI model within the AR context, the expert groups outlined how belief supports the filtering of duplicated or unnecessary assistance and acts as a screening step to narrow the assistance’s scope. They further outlined that desire refers to the goals of a given task. In AR, this means the system should model the user’s actions and goals (Group B and C). Finally, the expert groups indicated that intention comprises a small step toward the goal and affects the timing and content of the AR assistance (Groups A and C). Together these adaptations of the BDI model help to construct a novel pipeline toward proactive AR assistance.
@@
Table 3. The table summarizes the main results from two formative studies.
@@ 5. Design Requirements
    Assistance should try to be transparent to gain users’ trust, feed back the system’s reasoning and detection, and provide easily accessible information about current and overall goals in the AR environment.
@@ 6.1. Implementing the BDI Model for AR Assistance
Table 4. This table illustrates how three components— Belief, Desire, and Intention– in the BDI model are adapted for AR task guidance. BDI Comp. refers to the BDI components and AR Guid. Comp. refers to the AR system’s task guidance components. Belief is represented through scene understanding, task-relevant object detection, and user action history to minimize distractions, locate objects, and avoid repeated instructions. Desire captures the user’s high-level task goals, inferred through LLM-based scene analysis and confirmed by the user to ensure accuracy. Intention includes predicting the following intended action using GPT-based inference with chain-of-thought reasoning and determining the timing of next actions with checkpoint-based early forecasting.
@@
Figure 3. The figure is a system overview of the BDI user model. The system processes inputs from the camera’s view, dialogue (voice communication between the user and the GPT model), and the historical logger (records of prior assistance). These inputs are sent to different BDI components for analysis and inference using a combination of local models and LLMs to generate proactive guidance and determine the appropriate modality and assistance timing. To ensure assistance appears and disappears at the right time, a task planner LLM generates a step-by-step task plan based on the inferred desire, with multiple checkpoints assigned to each step. These checkpoints are monitored by the action finish detection module, which determines task completion by verifying checkpoint progress. In addition, the system employs an early forecasting module to minimize latency.
@@ 6.5. Task Planner for Checkpoints
Figure 4. Comparison of the naively generated images from the GPT model (i.e., Naive) with our proposed prompts (i.e., Satori). (a) “One hand presses a white button on a white espresso machine. A large red arrow points to the button. No background, in the style of flat, instructional illustrations. Accurate, concise, comfortable color style.” (b) “One hand presses a white button on a white espresso machine.” (c) “Cut stem of a red flower up from bottom, with white scissors at 45 degrees. One big red arrow pointing to bottom of the flower stem. In the style of flat, instructional illustrations. No background. Accurate, concise, comfortable color style.” (d) “Cut stem of a red flower up from bottom with white scissors at 45 degrees.
@@ 7.1. Conditions
Figure 6. Evaluation tasks using either Satori or a Wizard-of-Oz baseline. (a) The participant is assembling a mop during the room-cleaning task; and (b) The participant is connecting an HDMI cable to a Nintendo Switch dock during the connecting Nintendo Switch task; and (c) The participant is preparing a filter during the coffee-making task; and (d) The participant is trimming flower stems during the flower-arranging task.
@@ 8.1. System Evaluation
Satori 100.00 100.00
[…]
@@
Task Scene Understanding Object Understanding
HCRN 69.50 N/A
@@
[…]
L.A. Time GTEA Our Dataset
Recall Prec. F1 Recall Prec. F1
@@
[…]
@@ 8.2. Usability Scale
Table 9. The table summarizes the mean scores and 95% confidence intervals (CI) for each system (our Satori system and WoZ designed by the AR designer) across usability scale questions using non-inferiority tests. The “Vanilla” columns provide the Wilcoxon signed-rank test results (W statistic and p-values) for significant differences between systems. The “Non-Inferiority” columns show W statistics and p-values testing if Satori’s performance is non-inferior to WoZ within a set margin. The highlighted cells indicate established non-inferiority, suggesting that Satori performs comparably or better than WoZ over system performance and usability.
Figure 7. Color-coded seven-point Likert scale ratings are shown in the figure for the twelve-participant study. The figure compares the responses for Satori and WoZ systems across four tasks: Arranging Flowers, Making Coffee, Cleaning the Room, and Connecting a Console. Each bar represents the distribution of responses for a specific usability question, highlighting differences in user satisfaction, comprehensibility, and task support provided by both systems.
@@
Table 10. This table shows the results for NASA TLX questions and non-inferiority tests using the mean scores and 95% confidence intervals (CI) for Satori and WoZ systems across six dimensions: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. The Vanilla Wilcoxon signed-rank test results and non-inferiority tests (highlighted in green) indicate whether the Satori system performs comparably or better than the WoZ system in terms of cognitive load.
Figure 8. The box plot of NASA-TLX results illustrates the distribution of cognitive load ratings across six dimensions: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Each box represents the interquartile range (IQR) with the median marked by a horizontal line, showing the variability and central tendency of participants’ workload ratings for both systems. The comparison highlights differences in perceived workload between the WoZ and Satori conditions, providing insights into the effectiveness and usability of each approach.